Write a quadratic function f whose only zero is 13. f(x) = 0

never mind i got it
Answer #1

zero =13 Degree =2 so f(x) = (x-13)² foc) = +132-2 2 (13) fee) = x +169 -26% f(x) = x²- 268 +169 if(x)=x²- 26x +169)